OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 132= 1/2x+2(x-9) What is the regular price of a child's ticket?

OpenStudy (campbell_st):

solve the equation for x 1st expand the brackets and collect like terms 1/2x + 2x - 18 = 132 5/2 x - 18 = 132 add 18 to both sides 5/2x = 150 5x = 300 x = 60 the adult price is $60 the childs price is 60 - 9 = $51